A car moving with a speed of 180km/h was brought uniformly to rest by the application of the brakes in 20s.How far did the DA travel after the brakes were applied?
1
Expert's answer
2020-10-25T18:28:00-0400

"V_0= 180\frac{km}{h}=180*\frac{1000}{3600}\frac{m}{s}=50\frac{m}{s}"

"V= 0;\ a =\frac{V-V_0}{t}=\frac{0-50}{20}=-2.5\frac{m}{s^2}"

"S=Vt+\frac{at^2}{2}=50*20-\frac{2.5*20*20}{2}=500\ m"

Answer: 500 m